# Mathematical Theorems on Turbulence

### v1

@@ 5. Landau’s objection and intermittency
Where is the weak point in the argument we sketch for the behavior (5.1)? As we have seen from Theorem 4.3, the application of the $\nicefrac{{4}}{{5}}$ law is quite justified. This theorem also, in part, justifies the constancy of the flux in the inertial range since $\langle\varepsilon\rangle$ directly appears there. Well then, what is the evidence towards this non-intermittent dissipation postulate? As it turns out, this seems to be the biggest gap in Kolmogorov’s picture since in the typically 3D turbulence situation, it appears badly violated. Lev Landau was apparently the first to recognize thisOnsager also knew of intermittency. In his 1945 letter to C.C. Lin stated that energy spectral exponent might more negative than $\nicefrac{{5}}{{3}}$ because of “spottiness” of velocity-increments in space. [99], and point it out to Kolmogorov after a seminar in Kazan in 1942 (see discussion in Frisch [103] for a deeper historical analysis). In their text, Landau and Lifshitz [140].
@@
In the above, we have introduce a mollification so as to make use of the exact identity (5.2) in a more precise fashion. As a result, the mollified energy dissipation makes its way into the picture (this is also the basis for Kolmogorov’s own refinement of his theory from 1962 [132], which features ball averaged dissipation). See also Kraichnan [138]. At this point, Kolmgorov’s original argument relates the $\nicefrac{{p}}{{3}}$ moment of the dissipation simply to its mean. Instead, following Landau’s suggestion, a reasonable implication of the observed spottiness of the dissipation would be power-law divergences of higher moments
@@
Bounded weak solutions of Euler on $\mathbb{T}^{d}\times(0,T)$ with non-trivial energy dissipation measure ${\varepsilon}[u]$ have
@@
The proof is given in [59]. We prove the result here for the upper box counting dimension instead of Hausdorff, so that we do not need to appeal to Frostman’s lemma. Denote by ${\rm supp}\ \!{\varepsilon}]_{\delta}$ the space-time $\delta$-neighbourhood of ${\rm supp}\ \!{\varepsilon}$. Let $\chi^{\delta}\in C^{\infty}_{x,t}$ be so
@@
On $\mathbb{T}^{d}\times(0,T)$, let $\{u^{\nu}\}_{\nu>0}$ be a sequence of smooth solutions to Navier-Stokes. Assume that, in the limit as $\nu\rightarrow 0$, ${\varepsilon}^{\nu}[u^{\nu}]:=\nu|\nabla u^{\nu}|^{2}$ converges in $\mathcal{D}^{\prime}_{x,t}$ to a measure ${\varepsilon}[u]$ whose singular part with respect to the Lebesgue measure is non-trivial and concentrated on a set $S$ with $\dim_{\mathcal{H}}S=\gamma\in[1,d+1]$. For all $p\in[3,\infty]$ for which there exists $\zeta_{p}\in(0,p)$ such that $\{u^{\nu}\}_{\nu>0}$ stays bounded in $L^{p}_{t}B^{\frac{\zeta_{p}}{p}}_{p,\infty}$, it must hold that $S_{p}(\ell)\lesssim\ell^{\zeta_{p}}$ with
@@
One can easily check the above condition reduces to that reported on $\zeta_{p}$ above. To prove our claim, let $\varphi\in C^{\infty}_{x,t}$ be any compactly supported test function. Denote by ${\rm supp}\ \!{\varepsilon}]_{\delta}$ the space-time $\delta$-neighborhood of ${\rm supp}\ \!{\varepsilon}$. Let $\chi^{\delta}\in C^{\infty}_{x,t}$ be such that (5.7) holds. ThenFor simplicity us assume that there is no Lebesgue part of the dissipation measure. Indeed, if there were we could split ${\varepsilon}[u]={\varepsilon}_{\rm sing}[u]+{\varepsilon}_{\rm leb}[u]$. Our argument would apply to $\langle{\varepsilon}_{\rm sing}[u],\varphi\rangle=\langle{\varepsilon}_{\rm sing}[u],\varphi\chi_{\delta}\rangle=\langle{\varepsilon}[u],\varphi\chi_{\delta}\rangle-\langle{\varepsilon}_{\rm leb}[u],\varphi\chi_{\delta}\rangle$. For the first term, we proceed as in the proof. For the second it suffices to say that it tends to zero as $\delta\to 0$, as ${\varepsilon}_{\rm leb}[u]$ has an $L^{1}$ density and the set $S$ has dimension strictly less that $d+1$ (otherwise the statement is trivial). $\left\langle{\varepsilon}[u],\varphi\right\rangle=\left\langle{\varepsilon}[u],\varphi\chi^{\delta}\right\rangle$, and by the decomposition (3.21) we infer
@@
Since $\sigma,p,\gamma$ satisfy an open condition, we find $\alpha>0$ sufficiently small so that the exponent of $\delta$ in the above inequality is positive. The proof follow by contradiction of the non-triviality of ${\varepsilon}[u]$ by letting $\delta\rightarrow 0$. ∎
@@
Figure 14. Structure function exponents for $p\in[3,6]$ [60]. Blue dots are absolute structure function exponents measured from the JHU turbulence database. Red triangles are transverse exponents reported in [125]. Dashed black line corresponds to the Kolmogorov prediction of $\frac{p}{3}$. Solid grey line corresponds to $\zeta_{p}^{*}$ with $\gamma=3.85$.
@@
On $\mathbb{T}^{d}\times(0,T)$, let $u$ be is weak Euler solution with anomalous dissipation measure ${\varepsilon}[u]$. Assume $u\in L^{p}_{t}B^{\sigma_{p}}_{p,\infty}$ for some $p\in[3,\infty]$ and $\sigma_{p}\in\left(0,1\right)$. Then there exists and $\ell_{0}$ such that for all $\ell<\ell_{0}$
${\varepsilon}[u](B_{\ell}(x,t))\lesssim\ell^{\frac{2\sigma_{p}}{1-\sigma_{p}}-1+\frac{p-3}{p}(d+1)}\qquad\forall(x,t)\in\mathbb{T}^{d}\times(0,T).$ (5.9)
@@
Let us conclude the section with some discussion of the results. All the results of this section assume the dissipation measure charges some lower dimensional set, and derives implications of this to the structure of the velocity field. Specifically, we show the velocity cannot be too regular for this to happen, and the regularity cannot be uniform in $p$, at least for $p>3$. This implies that some form of intermittency must occur. On the other hand, we do not have any way of deducing that the dissipation measure must charge lower dimensional sets from measurements of the velocity field itself. One must instead measure the dissipation field directly, as was done by Meneveau and Sreenivasan. However, the multifractal model picture for Frisch and Parisi [106] (see also lecture notes of Eyink [83] and review of DeBrulle [54]) predicts that the slope of the curve of $\zeta_{p}$ at $p=0$, namely $h_{*}:=\frac{{\rm d}\zeta_{p}}{{\rm d}p}\Big|_{p=0}$ relates to the "most probable" Hölder exponent in the flow. Namely the one occupying a full space-time volume, with other local exponents $h\neq h_{*}$ occupying fractal sets of lesser dimension. Since $\zeta_{p}$ is a concave function of $p$, if any value of $\zeta_{p}>p/3$ for $p\in(0,3)$, then it follows that this exponent $h_{*}>1/3$. Thus, in light of Onsager’s theorem about $\nicefrac{{1}}{{3}}$ being the requisite singularity for dissipativity, one might guess that ${\varepsilon}[u]$ will charge lower dimensional if $\zeta_{p}>p/3$ for $p\in(0,3)$. We therefore end with the following
@@ 6. Lessons from model problems
6.1. Burgers’s equation for pressureless gas
@@
Let $\varphi:\mathbb{R}\to\mathbb{R}$ have Lipschitz first derivative. Let $\Phi$ be the primitive of $\varphi$ and $\Phi$} of $x\varphi^{\prime}(x)$. Then
@@
$J_{\ell}[u]$ $:=u^{\prime}\varphi(\delta_{\ell}u)-{\Phi}(\delta_{\ell}u)-\nu\partial_{x}\varphi(\delta_{\ell}u),$
$\Pi[\delta_{\ell}u]$ $:={\Phi}(\delta_{\ell}u)-\Phi(\delta_{\ell}u).$
@@
For $p\geq 3$ and $\alpha\in\mathbb{R}$, let $\varphi(x)=\alpha x|x|^{p-2}$. We compute $\varphi^{\prime}(x)=\alpha(p-1)|x|^{p-2}$ and $\varphi^{\prime\prime}(x)=\alpha(p-1)(p-2)x|x|^{p-4}\in L^{\infty}$. The primitive of $\varphi$ is $\Phi=\frac{\alpha}{p}|x|^{p}$ and for $x\varphi^{\prime}(x)=(p-1)\varphi(x)$, it is ${\Phi}=(p-1)\Phi$. Thus $\Pi[x]=(p-2)\Phi(x)=\alpha\frac{p-2}{p}|x|^{p}$. Letting $\alpha=\frac{p}{p-2}$, we obtain the balance
@@ 6.2. Obukhov, Corrsin, and Kraichnan’s scalar turbulence
Next we discuss passive advection and diffusion of scalar fields. Let $M\subset\mathbb{R}^{d}$ be open. Given a divergence-free vector field $u:\Omega\times(0,T)\rightarrow\mathbb{R}^{d}$, consider the advection-diffusion equation governing the motion of a tracer dye or temperature
@@
$\int_{0}^{T}\mathbb{E}\|\theta_{t}\|_{B^{1-\alpha}_{2,\infty}}^{2}{\rm d}t<\infty$ (6.23)
with the Besov-like spaces introduced in Lemma 2.1. In view of Theorem 6.8 and (6.22), the solution has sharply $1-\alpha$ derivatives in this $L^{2}$ based Besov space, at least in the mean sense, thereby establishing the analogue of Conjecture 4.5 for the Kraichnan model. This kind of regularization was established first in the work [108] by different means, and it shows the sharpness of the Obukhov-Corrsin regularity theory at the level of $L^{2}$, while allowing for intermittency [138]. In fact, it is remarkable that in the Kraichnan model model, an asymptotic approximation of the structure function scaling exponents $S_{p}^{\theta}(\ell):=\mathbb{E}|\delta_{\ell}\theta|^{p}\sim\ell^{\zeta_{p}^{\theta}}$ have been computed [19]
@@
$\frac{{\rm d}}{{\rm d}t}\fint_{B_{\ell}}S^{\kappa}_{t}(z){\rm d}z=\fint_{B_{\ell}}\nabla\cdot(Q\nabla S^{\kappa}_{t})(z){\rm d}z-2\kappa\fint_{B_{\ell}}\mathbb{E}[\|\nabla\delta_{z}\theta^{\kappa}_{t}\|_{L^{2}_{x}}^{2}]{\rm d}z$
@@ 7. Taylor, Richardson and the motion of particles
Should we expect this phenomenon in real turbulence? One of the first investigations on particles motion in turbulence is due to G.I. Taylor [182] (1921), who studied how a single particle diffuses at long time. Taylor predicted that a single particle acts diffusively at long times with effective diffusivity $D$, e.g.
@@
provided the latter limit exists. Here $\langle\cdot\rangle$ denotes average over the label $a\in\mathbb{T}^{d}$, and possibly over an ensemble of velocity fields. Taylor’s result holds under the assumption that the Lagrangian velocity $v_{t}(a)=\dot{X}_{t}(a)$ is stationary in time after averaging, namely $\langle v_{t}(a)\cdot v_{t^{\prime}}(a)\rangle=\langle v_{t-t^{\prime}}(a)\cdot v_{0}(a)\rangle$, which should be an integrable function of $\tau=t-t^{\prime}$. Stationarity of the Lagrangian velocity holds, for instance, provided the Eulerian velocity field is statistical stationary and homogeneous. See [183, §7.1] and [133] for an example of an early rigorous statement. Thus, we take it that a single particle looks like Brownian motion at long times, albeit with a diffusivity which is only implicitly predicted.
@@
Of course, there are simple examples saturating this upper bound (take $u(x)={\rm sgn}(x)|x|^{\alpha}$ in one dimension), showing the Proposition to be sharp. See [93, 92] for a realization of this simple example in quantum systems, and a renormalization group perspective. The works [75, 74] give higher dimensional examples and investigates when a unique measure valued solution might be obtained. The works [149, 150, 151, 184] concern the same behavior for infinite dimensional models, where these features robustly appear.
@@
Although Richardson’s prediction has proved difficult to conclusively observe, impressive partial confirmations have been made experimentally [162, 23] and numerically [21, 24, 17, 22]. The following Figure 24, taken from [22], represents essentially the state of the art in our observations of Richardson’s law from numerical simulation data. The $t^{3}$ law is observed, as is the "forgetting" of the initial separations as expected from Richardson’s theory as well as the simple-minded bound in Proposition 7.1.
@@
It is unfortunate that many of the rigorous results discussed in this section require Hölder continuity as opposed to Besov. For instance, given the observations do not support $\nicefrac{{1}}{{3}}$ Holder continuity, we wonder if there is an analogous Besov statement for Lagrangian regularity. Namely, if $u\in L_{t}^{p}(0,T;B_{p,\infty}^{\sigma}(\mathbb{T}^{d}))$, is it true that $v=L^{p/2}(\mathbb{T}^{d};B^{\frac{\sigma}{1-\sigma}}_{p/2,\infty}(0,T))$? The difficulty in establishing this type of result is that, unlike the Hölder case, it seems necessary to bundle trajectories and the regularity in the label $a$ can be very bad. It may be that is impossible to bundle trajectories in such as way as to define a measure-preserving flowmap $X_{t}(a)$ that depends measurably on $a$ (see [164] for an example). Could such regularity at least be preserved uniformly along solutions of Navier-Stokes?

### v2

@@ 5. Landau’s objection and intermittency
Where is the weak point in the argument we sketch for the behavior (5.1)? As we have seen from Theorem 4.3, the application of the $\nicefrac{{4}}{{5}}$ law is quite justified. This theorem also, in part, justifies the constancy of the flux in the inertial range since $\langle\varepsilon\rangle$ directly appears there. Well then, what is the evidence towards this non-intermittent dissipation postulate? As it turns out, this seems to be the biggest gap in Kolmogorov’s picture since in the typically 3D turbulence situation, it appears badly violated. Lev Landau was apparently the first to recognize thisOnsager also knew of intermittency. In his 1945 letter to C.C. Lin stated that energy spectral exponent might more negative than $\nicefrac{{5}}{{3}}$ because of “spottiness” of velocity-increments in space. [99], and pointed it out to Kolmogorov after a seminar in Kazan in 1942 (see discussion in Frisch [103] for a deeper historical analysis). In their text, Landau and Lifshitz [141].
@@
In the above, we have introduced a mollification so as to make use of the exact identity (5.2) in a more precise fashion. As a result, the mollified energy dissipation makes its way into the picture (this is also the basis for Kolmogorov’s own refinement of his theory from 1962 [133], which features ball averaged dissipation). See also Kraichnan [139]. At this point, Kolmgorov’s original argument relates the $\nicefrac{{p}}{{3}}$ moment of the dissipation simply to its mean. Instead, following Landau’s suggestion, a reasonable implication of the observed spottiness of the dissipation would be power-law divergences of higher moments
@@
Bounded weak solutions of Euler on $(0,T)\times\mathbb{T}^{d}$ with non-trivial energy dissipation measure ${\varepsilon}[u]$ have
@@
The proof is given in [59]. We prove the result here for the upper box counting dimension instead of Hausdorff, so that we do not need to appeal to Frostman’s lemma. Denote by ${\rm supp}\ \!{\varepsilon}]_{\delta}$ the space-time $\delta$-neighbourhood of ${\rm supp}\ \!{\varepsilon}$. Let $\chi^{\delta}\in C^{\infty}_{x,t}$ be such that
@@
On $(0,T)\times\mathbb{T}^{d}$, let $\{u^{\nu}\}_{\nu>0}$ be a sequence of smooth solutions to Navier-Stokes. Assume that, in the limit as $\nu\rightarrow 0$, ${\varepsilon}^{\nu}[u^{\nu}]:=\nu|\nabla u^{\nu}|^{2}$ converges in $\mathcal{D}^{\prime}_{x,t}$ to a measure ${\varepsilon}[u]$ whose singular part with respect to the Lebesgue measure is non-trivial and concentrated on a set $S$ with $\dim_{\mathcal{H}}S=\gamma\in[1,d+1]$. For all $p\in[3,\infty]$ for which there exists $\zeta_{p}\in(0,p)$ such that $\{u^{\nu}\}_{\nu>0}$ stays bounded in $L^{p}_{t}B^{\frac{\zeta_{p}}{p}}_{p,\infty}$, it must hold that $S_{p}(\ell)\lesssim\ell^{\zeta_{p}}$ with
@@
One can easily check the above condition reduces to that reported on $\zeta_{p}$ above. To prove our claim, let $\varphi\in C^{\infty}_{x,t}$ be any compactly supported test function. Denote by ${\rm supp}\ \!{\varepsilon}]_{\delta}$ the space-time $\delta$-neighborhood of ${\rm supp}\ \!{\varepsilon}$. Let $\chi^{\delta}\in C^{\infty}_{x,t}$ be such that (5.7) holds. ThenFor simplicity, let us assume that there is no Lebesgue part of the dissipation measure. Indeed, if there were we could split ${\varepsilon}[u]={\varepsilon}_{\rm sing}[u]+{\varepsilon}_{\rm leb}[u]$. Our argument would apply to $\langle{\varepsilon}_{\rm sing}[u],\varphi\rangle=\langle{\varepsilon}_{\rm sing}[u],\varphi\chi_{\delta}\rangle=\langle{\varepsilon}[u],\varphi\chi_{\delta}\rangle-\langle{\varepsilon}_{\rm leb}[u],\varphi\chi_{\delta}\rangle$. For the first term, we proceed as in the proof. For the second it suffices to say that it tends to zero as $\delta\to 0$, as ${\varepsilon}_{\rm leb}[u]$ has an $L^{1}$ density and the set $S$ has dimension strictly less that $d+1$ (otherwise the statement is trivial). $\left\langle{\varepsilon}[u],\varphi\right\rangle=\left\langle{\varepsilon}[u],\varphi\chi^{\delta}\right\rangle$, and by the decomposition (3.21) we infer
@@
Since $\sigma,p,\gamma$ satisfy an open condition, we find $\alpha>0$ sufficiently small so that the exponent of $\delta$ in the above inequality is positive. The proof follows by contradiction of the non-triviality of ${\varepsilon}[u]$ by letting $\delta\rightarrow 0$. ∎
@@
Figure 14. Structure function exponents for $p\in[3,6]$ [60]. Blue dots are absolute structure function exponents measured from the JHU turbulence database [129]. Red triangles are transverse exponents reported in [125]. Dashed black line corresponds to the Kolmogorov prediction of $\frac{p}{3}$. Solid grey line corresponds to $\zeta_{p}^{*}$ with $\gamma=3.85$.
@@
On $(0,T)\times\mathbb{T}^{d}$, let $u$ be a weak Euler solution with anomalous dissipation measure ${\varepsilon}[u]$. Assume $u\in L^{p}_{t}B^{\sigma_{p}}_{p,\infty}$ for some $p\in[3,\infty]$ and $\sigma_{p}\in\left(0,1\right)$. Then there exists an $\ell_{0}$ such that for all $\ell<\ell_{0}$
${\varepsilon}[u](B_{\ell}(x,t))\lesssim\ell^{\frac{2\sigma_{p}}{1-\sigma_{p}}-1+\frac{p-3}{p}(d+1)}\qquad\forall(x,t)\in(0,T)\times\mathbb{T}^{d}.$ (5.9)
@@
Let us conclude the section with some discussion of the results. All the results of this section assume the dissipation measure charges some lower dimensional set, and derives implications of this to the structure of the velocity field. Specifically, we show the velocity cannot be too regular for this to happen, and the regularity cannot be uniform in $p$, at least for $p>3$. This implies that some form of intermittency must occur. On the other hand, we do not have any way of deducing that the dissipation measure must charge lower dimensional sets from measurements of the velocity field itself. One must instead measure the dissipation field directly, as was done by Meneveau and Sreenivasan. However, the multifractal model picture for Frisch and Parisi [106] (see also lecture notes of Eyink [83] and review of DuBrulle [54]) predicts that the slope of the curve of $\zeta_{p}$ at $p=0$, namely $h_{*}:=\frac{{\rm d}\zeta_{p}}{{\rm d}p}\Big|_{p=0}$ relates to the "most probable" Hölder exponent in the flow. Namely the one occupying a full space-time volume, with other local exponents $h\neq h_{*}$ occupying fractal sets of lesser dimension. Since $\zeta_{p}$ is a concave function of $p$, if any value of $\zeta_{p}>p/3$ for $p\in(0,3)$, then it follows that this exponent $h_{*}>1/3$. Thus, in light of Onsager’s theorem about $\nicefrac{{1}}{{3}}$ being the requisite singularity for dissipativity, one might guess that ${\varepsilon}[u]$ will charge lower dimensional if $\zeta_{p}>p/3$ for $p\in(0,3)$. We therefore end with the following
@@ 6. Lessons from model problems
6.1. Burgers’ equation for pressureless gas
@@
Let $\varphi:\mathbb{R}\to\mathbb{R}$ have Lipschitz first derivative. Let $\Phi$ be the primitive of $\varphi$ and $\tilde{\Phi}$ of $x\varphi^{\prime}(x)$. Then
@@
$J_{\ell}[u]$ $:=u^{\prime}\varphi(\delta_{\ell}u)-\tilde{\Phi}(\delta_{\ell}u)-\nu\partial_{x}\varphi(\delta_{\ell}u),$
$\Pi[\delta_{\ell}u]$ $:=\tilde{\Phi}(\delta_{\ell}u)-\Phi(\delta_{\ell}u).$
@@
For $p\geq 3$ and $\alpha\in\mathbb{R}$, let $\varphi(x)=\alpha x|x|^{p-2}$. We compute $\varphi^{\prime}(x)=\alpha(p-1)|x|^{p-2}$ and $\varphi^{\prime\prime}(x)=\alpha(p-1)(p-2)x|x|^{p-4}\in L^{\infty}$. The primitive of $\varphi$ is $\Phi=\frac{\alpha}{p}|x|^{p}$ and for $x\varphi^{\prime}(x)=(p-1)\varphi(x)$, it is $\tilde{\Phi}=(p-1)\Phi$. Thus $\Pi[x]=(p-2)\Phi(x)=\alpha\frac{p-2}{p}|x|^{p}$. Letting $\alpha=\frac{p}{p-2}$, we obtain the balance
@@ 6.2. Obukhov, Corrsin, and Kraichnan’s scalar turbulence
Next we discuss passive advection and diffusion of scalar fields. Let $M\subset\mathbb{R}^{d}$ be open. Given a divergence-free vector field $u:(0,T)\times M\rightarrow\mathbb{R}^{d}$, consider the advection-diffusion equation governing the motion of a tracer dye or temperature
@@
$\int_{0}^{T}\mathbb{E}\|\theta_{t}\|_{B^{1-\alpha}_{2,\infty}}^{2}{\rm d}t<\infty,$ (6.23)
with the Besov-like spaces introduced in Lemma 2.1. In view of Theorem 6.8 and (6.22), the solution has sharply $1-\alpha$ derivatives in this $L^{2}$ based Besov space, at least in the mean sense, thereby establishing the analogue of Conjecture 4.5 for the Kraichnan model. This kind of regularization was established first in the work [108] by different means, and it shows the sharpness of the Obukhov-Corrsin regularity theory at the level of $L^{2}$, while allowing for intermittency [139]. In fact, it is remarkable that in the Kraichnan model, an asymptotic approximation of the structure function scaling exponents $S_{p}^{\theta}(\ell):=\mathbb{E}|\delta_{\ell}\theta|^{p}\sim\ell^{\zeta_{p}^{\theta}}$ have been computed [19]
@@
$\frac{{\rm d}}{{\rm d}t}\fint_{B_{\ell}}S^{\kappa}_{t}(z){\rm d}z=\fint_{B_{\ell}}\nabla\cdot(Q\nabla S^{\kappa}_{t})(z){\rm d}z-2\kappa\fint_{B_{\ell}}\mathbb{E}[\|\nabla\delta_{z}\theta^{\kappa}_{t}\|_{L^{2}_{x}}^{2}]{\rm d}z.$
@@ 7. Taylor, Richardson and the motion of particles
Should we expect this phenomenon in real turbulence? One of the first investigations on particles motion in turbulence is due to G.I. Taylor [183] (1921), who studied how a single particle diffuses at long time. Taylor predicted that a single particle acts diffusively at long times with effective diffusivity $D$, e.g. with the Lagrangian velocity $v_{t}(a):=\dot{X}_{t}(a)$,
@@
provided the latter limit exists. Here $\langle\cdot\rangle$ denotes average over the label $a\in\mathbb{T}^{d}$, and possibly over an ensemble of velocity fields. Taylor’s result holds under the assumption that the Lagrangian velocity is stationary in time after averaging, namely $\langle v_{t}(a)\cdot v_{t^{\prime}}(a)\rangle=\langle v_{t-t^{\prime}}(a)\cdot v_{0}(a)\rangle$, which should be an integrable function of $\tau=t-t^{\prime}$. Stationarity of the Lagrangian velocity holds, for instance, provided the Eulerian velocity field is statistical stationary and homogeneous. See [184, §7.1] and [134] for an example of an early rigorous statement. Thus, we take it that a single particle looks like Brownian motion at long times, albeit with a diffusivity which is only implicitly predicted.
@@
Of course, there are simple examples saturating this upper bound (take $u(x)={\rm sgn}(x)|x|^{\alpha}$ in one dimension), showing Proposition 7.1 to be sharp. See [93, 92] for a realization of this simple example in quantum systems, and a renormalization group perspective. The works [75, 74] give higher dimensional examples and investigates when a unique measure valued solution might be obtained. The works [150, 151, 152, 185] concern the same behavior for infinite dimensional models, where these features robustly appear.
@@
Although Richardson’s prediction has proved difficult to conclusively observe, impressive partial confirmations have been made experimentally [163, 23] and numerically [21, 24, 17, 22]. The following Figure 24, taken from [22], represents essentially the state of the art in our observations of Richardson’s law from numerical simulation data. The $t^{3}$–law is observed, as is the "forgetting" of the initial separations as expected from Richardson’s theory as well as the simple-minded bound in Proposition 7.1.
@@
It is unfortunate that many of the rigorous results discussed in this section require Hölder continuity as opposed to Besov. For instance, given the observations do not support $\nicefrac{{1}}{{3}}$ Hölder continuity, we wonder if there is an analogous Besov statement for Lagrangian regularity. Namely, if $u\in L_{t}^{p}(0,T;B_{p,\infty}^{\sigma}(\mathbb{T}^{d}))$, is it true that $v=L^{p/2}(\mathbb{T}^{d};B^{\frac{\sigma}{1-\sigma}}_{p/2,\infty}(0,T))$? The difficulty in establishing this type of result is that, unlike the Hölder case, it seems necessary to bundle trajectories and the regularity in the label $a$ can be very bad. It may be that is impossible to bundle trajectories in such as way as to define a measure-preserving flowmap $X_{t}(a)$ that depends measurably on $a$ (see [165] for an example). Could such regularity at least be preserved uniformly along solutions of Navier-Stokes?